Grade 11Physical Chemistry

an automobile antifreeze consists of 38.7 % c9.7 % h and remaining oxygen by weight when 0.93 g of it are vaporized at STP 336 ml of vapour are formed find the molecular formula of the anti freeze

Gaurang Kotasthane
10 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago
element : C H O
% of mass : 38.7 9.7 51.6
mass of element
in 100g of compound: 38.7g 9.7g 51.6g
mole : 38.7/12 9.7/1 51.6/16
=3.22 =9.7 =3.22
mole of ratio : 3.22/3.22 9.7/3.22 3.22/3.22
=1(2) =3(2) =1(2)
E.F= C2H6O2